OpenStudy (anonymous):

PLEASE HELP!! :( Derivative of r = tan(x)sec(x) a. = sec^2x(sin^3x+1) b. = sec^3x(sin^2x +1) c. = sin^3 x (sec^2x +1) d. = sec^3x (sinx+1)

OpenStudy (fibonaccichick666):

so, are you familiar with the product rule?

OpenStudy (anonymous):

\[\frac{ d }{dx }\left( uv \right)=uv'+vu'\]

OpenStudy (anonymous):

yes! i know the product rule.

OpenStudy (fibonaccichick666):

so, what you are going to do is apply it, then simplify

OpenStudy (fibonaccichick666):

so why don't you try the first step

OpenStudy (anonymous):

\[\frac{ d }{ dx }\tan x=\sec ^2x\] \[\frac{ d }{ dx }\left( \sec x \right)=\sec x~ \tan x\]

OpenStudy (anonymous):

what i got when i used the product rule was sec(x)(tan^2(x)+sec^2(x))

OpenStudy (anonymous):

but i could not figure out how the sin(x) got into the equation...

OpenStudy (fibonaccichick666):

ok good, so do you know any identities that use tan^2x or sec^2 x?